A police van moving on a highway with a speed of 30km/hr fires a bullet at a thiefs car speeding away in the same direction with a speed of 192km/hr.if the muzzle speed of the bullet is 150m/s with what speed does the bullet hit the thiefs car??

Grade:

5 Answers

amit phogat
10 years ago

105m/s

SOURAV MISHRA
10 years ago

first note that it is the speed of the bullet relative to the thief''s car with which the bullet hits it as seen by the thief. it is this speed and momentum which is responsible for the actual damage of the car.

speed of the bullet relative to the ground is 150 + 8.33 = 158.33 m/s.

as speed of bullet relative to ground = muzzle speed of bullet + speed of gun relative to ground.

the speed of the thief''s car relative to the ground is 53.33 m/s.

therefore the speed of the bullet relative to the thief''s car is 158.33 - 53.33 = 105 m/s. 

jakira
7 years ago
i  think km/r chnage m/s (5/18)..why u use 5/8 in this problem..please explain that concept …...................................
Munazir
6 years ago
First V ba , Bullet with respect to police van = 192-30 = 162 Then thief`s car w.r.t vba = 540 - 162 = 378 ,,, its ur answer ,, but convert 150 m/s to km/ h then 540 come
Roma
6 years ago
Velocity of police van( Vp) = 30 km\ h= 25/3 mpsVelocity of theif ( Vt) = 192 kph = 160/3 mpsMuzzle velocity ( Vbg) = Vb - Vg = 150 mps Vg = Vp = 25/3 mpsVb = 475/3 mpsDamaging speed = velocity of bullet with respect to thief car = Vbt = Vb - Vt = 475/3 - 160 /3 ` = 105 mps